OpenStudy (anonymous):

Solve each system by elimination: -5x+9y=-4 15x-4y=-11 please i need help

OpenStudy (anonymous):

multiply the first equation by 3. Then add down and the -15x will cancel with 15x elminating it! Leaving you able to solve for y. Back substitute into either equation to get x once you find y.

OpenStudy (anonymous):

I'm still confused....sorry...@morganashley can you help me?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

ok- so for solving by elimination, you need to add the equations together in a way that one variable will disappear. in this case, the x's look like they will be easier to eliminate first, because -5 can be multiplied by 3 to get 15. (-5x+9y=-4)*3 would give you -15x+36y=-12 +(15x- 4y=-11) ___________________ does that make sense? can you see that if you take the equations and add them, the 15x will be eliminated?

OpenStudy (anonymous):

sorry i was trying to finish typing it out. if you need more explanation let me know.

OpenStudy (anonymous):

sorry it still doesnt make sence i dont know where you got 3 from and what do you mean by adding the equations together,can you like draw it?

OpenStudy (anonymous):

ok. here we go. it will take a minute so hang on.

OpenStudy (anonymous):

  ok. so, to eliminate a variable, we need to have a positive and a negative coefficient of the variable. that is, we need a -15x and a +15x. the -15x is already there, so we need a -15x. we see in the first equation a -5x, which, if multiplied by 3, will give us the -15x we need. So, we multiply that full equation by 3 ( what ever you do to one thing, we have to do to everything in the equation). then we get -15x+27y=-12. (there was an error in my first answer. i apologize). now we add the equations together, like we would add any other numbers, like i tried to show in the second picture. add the x's, the y's and the numbers at the end. now the x's are gone, and we have 23y=-23. does this make sense so far?